### v2

@@ 4 Framework for LLM Advertisement
Prediction module is responsible for computing user’s satisfaction-rate (SR) and click-through-rate (CTR). User’s SR indicates how much the user is satisfied with the given output. This SR affects the final decision making process of LAS since if there is a chance that the user is very disappointed with the modified output, then it should reconsider displaying that output as it may hurt the user experience a lot. On the other hand, CTR indicates the probability that the user clicks the ad link included within the modified output. This is also critical in determining the auction winner since this directly affects the revenue of the LAS. More precisely, if an advertiser’s bidding method is cost-per-click (CPC), then CTR proportionally affects the advertisement revenue as the expected revenue of transferring its ad can be computed as CPC times the CTR. Overall, both of SR and CTR can be defined as a function of original output $X$, modified output $X^{\prime}$, query $q$, and context $c$, which returns a real value in $[0,1]$. Detailed methodologies for estimating/learning these functions will be discussed in Section 5.
Auction module
@@
Computed all the required parameters above, we finally run the auction module to determine who wins the auction and how much the advertiser will be charged. The input to the auction module is the tuples of $(\text{bid}_{i},\text{sr}_{i},\text{ctr}_{i})_{i\in[n]}$ which denotes bid amount, satisfaction rate, and click-through-rate, respectively. Output of the auction module is an allocation $a\in\{0,1\}^{n}$ and a payment $p\in\mathbb{R}_{\geq 0}$. Precisely, auction module determines the auction format which consists of the allocation function and the payment function, where the allocation function decides which advertisement to deliver and the payment function computes how much the advertisers pay LAS. [^9] In its current form, this does not allow the possibility of adjusting the final output in order to balance between multiple advertiser’s preferred output. Our framework, however, does not preclude possibility for such extension, which we will elaborate more in Appendix B.
@@
Importantly, the main goal of LAS is to maximize its revenue in a long term manner. To this end, this should balance a trade-off between short-term revenue and the user retention. Therefore, the objective can be modeled as a function from bid amount, CTR, and SR to a nonnegative real number that indicates a score of a modified output, i.e., selecting $i^{*}=\operatorname*{argmax}_{i\in[n]}\text{Obj}(\text{sr}_{i},\text{ctr}_{i},\text{bid}_{i})$. We do not discuss how to choose the objective function here, given the vast literature on the sponsored search auction. After designing the score function, one should determine an auction format. Also, Given the long line of literature on auction design, one can pick the most preferred mechanism, e.g., first-price auction (FPA) or second-price auction (SPA).
@@ 5 Challenges
Recall that our overall framework consists of four modules: modification, bidding, prediction, and auction module. For each module, we discuss several characteristics and technical challenges in implementing the proposed architectures in practice, evaluating their overall practicality with respect to the criteria defined in Section 3. Further discussions can be found in Appendix B due to page limit.
@@ 5.1 Modification module
Comparing modification models
In the advertiser modification model, LAS should somehow deliver $q,X,C$ to each $\text{adv}_{i}$. This may leak some private information of the user to the bidders, thereby suffering privacy issues, which is critical in terms of the user experience. Partially disclosing the information may help resolve the issue.
In addition, the advertiser modification model has reliability issue since there is a possibility that the advertiser-modified output includes some illegal or spamming contents. The LAS may devise an additional module to maintain its robustness against such kind of adversarial behavior, however, this is still a cost of the LAS to bring them in the system. This further possibly degrades the user satisfaction since the advertiser-generated output may not be satisfactory for the user compared to what the LLM generates. Furthermore, this requires additional communication between the LAS and the advertisers, thereby increasing the overall latency. Overall, advertiser modification model may not be feasible in practice due to several issues discussed above.
On the other hand, LAS modification model does not suffer the privacy, reliability and latency issues. Furthermore, this model would be more in favor of the user since LAS would basically hope to improve the user experience. However, one disadvantage is that this may not fully reflect the advertiser’s preference on the output modification process, thereby partly decreasing the advertiser’s satisfaction. This issue can be resolved by enabling more flexibility in the architecture, which will be discussed in what follows.
Beyond advertiser and LAS modification
In order to improve the advertiser’s satisfaction for the LAS modification module, one approach is to allow the advertiser to submit indirect indicators that represent its preference on what it expects the modified output to be. More in a detail, once the user asks a query $q$, LAS delivers $q$, the original output $X$, and possibly the context $c$ to the advertiser, and the advertiser returns a string $Y$ that indicates its preference in words. After then, LAS may apply $Y$ as a prompt in the LLM to generate the modified output. This approach allows more flexibility for the advertiser to express its preference on the output. Note that, however, this still requires additional communication cost compared to the LAS modification, which possibly increases latency.
@@ 5.2 Bidding module
Comparing bidding models
The biggest advantage of the dynamic bidding model is that, it would induce higher advertiser’s satisfaction rate, thereby attracting more advertisers to participate. This is because the advertisers have full flexibility in deciding their bids after observing the modified output. Similar to the advertiser modification model, dynamic bidding model may have privacy issue if LAS discloses some private information to the advertiser. Moreover, dynamic bidding model may suffer additional latency issue since the entire set of modified output should be delivered to the advertisers. Note that there is no reliability concern as only the bid amount is communicated.
Static bidding model, however, does not suffer any of the privacy, latency, reliability issues. However, as the advertisers cannot fully customize their bid per each query based on the modified output, this may degrade the advertiser satisfaction.
Beyond static and dynamic bidding
In the static bidding model, a natural approach to improve the advertiser’s satisfaction is to enable a more generic static function as the contract between the advertiser and LLM provider. For example, in a simplistic approach, the advertiser may specify specific user segments to deliver their advertisements, e.g., children under 16, or may screen out some traffics based on geographical information, e.g., users far from national parks. More critically, since the advertiser may want to provide a tailored bid with respect to the quality of its modified output, LAS can propose the bidding contract as a function between the quality of the modified output to a non-negative real number. For instance, the bidding function on the committed contract can be a linear function that is proportional to the distance measure between the modified output and the original output, e.g., charge \$$\$5\cdot d$ if the similarity distance is $d$. Note that in this case, the advertiser may need to partly understand how LAS estimates the distance measure and how it’s defined. More generally, LAS might construct more refined way to measure the units of interest, attention, relevance for the modified output in terms of $\text{adv}_{i}$, and let the advertisers commit to either of the contract based on its taste.
@@ 5.3 Prediction module
The process of estimating CTR will be analogous to that in the modern online advertisement system. We can train a prediction system to predict $\text{ctr}_{i}\in[0,1]$ given the input $X_{i},q,c$ for each $i\in[n]$ based on the feedback data from the user. [^10] In practice, one may extract some useful features from each $q,X$ and $c$, and learn the mapping from features to the true value. Feedback data will be the direct action of the user indicating whether the user clicks the ad or not. This prediction system can be trained in an online manner to update the estimate in (almost) real time, which will increase the accuracy of the prediction.
@@
To estimate SR, one may first consider an explicit approach given an assumption that the original output is optimal. If we assume that the original output is optimal, then essentially a distance between the original output $X$ and the modified output $X_{i}$ will be a good measure to predict the SR since if $X_{i}=X$, then the user will be the most satisfied. In this context, given a string $s(X)$ that represents the output $X$, we propose that the marginal probability ${\mathrm{Pr}\left[\,s(X)|q\,\right]}$ as an indicator for SR. Precisely, we define an output distance by $d(X,X_{i}):=d({\mathrm{Pr}\left[\,s(X)|q\,\right]},{\mathrm{Pr}\left[\,s(X_{i})\,\right]}|q)$ where $d$ denotes some valid metric. Note that the probability for an output to be obtained can be computed via standard method in the literature (e.g., see Vaswani et al. (2017)). Further, one can consider more general distance function $d$ to improve the accuracy of the prediction, such as functions to measure the semantic similarity of two phrases (Mikolov et al., 2013; Cer et al., 2018; Conneau et al., 2019).
@@
Alternatively, we can learn SR via feedback data from the user. A useful indicator of whether the user is satisfied with the output is whether the user re-generates the output given the same (or similar) query. Hence, one may aim to learn a function which outputs the sr, given the query $q$, modified output $X^{\prime}$, and context $c$.  [^11] For instance, as is typical in the modern online advertisement system, one may use factorization machine based online algorithm after extracting some features from user, context, and the queries. More detailed discussion can be found in McMahan et al. (2013). This approach does not assume that the original output $X$ is indeed optimal, thereby allowing the possibility that the user may be satisfied with a modified output $X_{i}$ even though its distance from $X$ is measured to be far. However, one should implement an independent training system to operate this learning process, leading to additional burden for LAS. To incorporate advantages of these two approaches, one can consider an online learning model where the distance is also integrated as one of the features to predict it, i.e., $d(X,X_{i})$ is used as a feature of the SR prediction system. In case $d(X,X_{i})$ indeed has some positive correlations with the true user satisfaction rate, this would increase the accuracy of the prediction.
@@ 5.4 Auction module
Multiple ads at once
Recall that in our auction module, the modified output is fully determined based on the allocation function so that if it is allocated to $\text{adv}_{i}$, then the modified output $X_{i}$ will be displayed to the user. One approach to generalize this framework is, to let the final output to be displayed the one that does not necessarily belong to $\{X_{i}\}_{i\in[n]}$, but rather a new output $X^{\prime}$ that somehow interpolates $\{X_{i}\}_{i\in[n]}$. [^12] We discuss further comparison to the token auction model by Duetting et al. (2023) in Appendix B.1. Let us denote the new output $X^{\prime}$ by balanced output. By doing so, it might be possible to deliver multiple advertisements in a fair manner, [^13] Another way to display multiple ads at once will is discussed in Appendix B.2 under our original framework. thereby allowing LAS to bring more revenue by charging multiple advertisers at once. One subtle issue is that, since each advertiser bids $b_{i}$ on delivering $X_{i}$, they may not want to bid the same amount for the balanced output $X^{\prime}$, thereby degrading the advertiser’s experience on participating in the LAS. In the static bidding model, as discussed in Section 5.2, this can be handled by committing to a contract based on measures that represents the advertiser’s preferences more in a refined manner. In the dynamic bidding model, of course, this may simply be handled by incorporating additional steps of querying the bid for the final output again.
@@ 6 Dynamic Creative Optimization via LLM
Figure 3: Integrating personalized advertisement in the images generated by ChatGPT 4. In the left panel, a user requests car recommendations within the price range of \$40,000 to \$60,000, and the chat bot, when capable of advertising, generates a “BMW 3 Series” image (left). Personalized ads modify this image for a 60-year-old woman (middle) or man (right) based on user context like age and gender. In the right panel, a user seeks machine learning conferences in Florida. The chat bot advertises “SunShine Travel” with its logo on a generic Florida image (left). With personalized ads, the model tailors promotions, offering flight tickets from the user’s location and showing more relevant Florida images. For example, for a young unmarried person from DC, the output advertises DC to Florida flights (middle) and features an image of beach there, while for a married person in California, it promotes California to Florida flights with images of amusement parks. This enhances user experience with targeted and appealing content.
@@
Now, we discuss about the possibility of advanced dynamic creative optimization by using LLM. Dynamic creative optimization (DCO)  [^14] This is often called as dynamic product ads (DPA), but DPA typically refers to the process of selecting which item to advertise from a given catalog of an advertiser, whereas DCO denotes the process of modifying the ad itself., in the traditional search and display ad markets, refers to the technique of dynamically adjusting the combination of ad assets and constructing an advertisement that best fits the customer’s preference. It can markedly enhance the quality and relevance of the ad to the user by tailoring the contents of ad in a personalized manner. Many online advertisement platforms adopt DCO-based ads, which is often called as responsive ad GoogleDisplayAd ; GoogleSearchAd .  [^15] As traditional DCO is restricted to dynamically construct the complete image file of ad, we here restrict our attention to the image-based ads as well. However, it is apparent that LLM based DCO would work beyond image-based ads.
@@
In the canonical DCO framework, a static ad consists of a single image file that can only be displayed in a specific set of ad channels, whereas dynamically optimized ad consists of a single ad template and multiple options for each asset that fills in the ad template. This is typically done by the advertiser registering multiple options for each asset given an ad template, and then the system picks the optimal combination. This procedure aims at increasing the CTR and SR, which would eventually improve the welfare efficiency of the online ad ecosystem McMahan et al. (2013) as well as the revenue of the ad platform.
@@
Unlike from the traditional DCO techniques, the arise of LLM, especially those support the output with image such as ChatGPT 4, enables more involved process of dynamic creation of ads, as well as replaces the role of traditional DCO in online ad platforms. For example, given a user query and context, the platform may send a query to LLM to customize the image of ad to attract the user by capturing user preference. We provide representative examples in Figure 3 where two scenarios of personalizing ads is depicted.
Incorporating responsive ad
@@
To integrate responsive ads, a more advanced modification module is imperative. This module should not only enable the alteration of the original response to include ads but also incorporate them in a manner that takes into account user preferences. This entails utilizing user contexts, e.g., gender/location/the current device used for the query, in the ad generation process. To be more specific, the modification module should incorporate user context in the process of modifying the original response. This makes ads more appealing to the user by considering their preferences. Leveraging language models, we can prompt them to factor in various elements related to user context during the generation of modified outputs. As illustrated in Figure 3, information about users’ context can be utilized to create more captivating advertisements for them. Further, the prediction module may interact with the modification module to commit to the modified output that mostly increases user’s experience, i.e., CTR and SR.
Cost sharing model
@@
Unlike from the traditional DCO technique that simply determines the efficient combination of ad assets, LLM-based responsive ad can do much beyond by creating truly new content. On the other hand, such process necessarily entails increased use of computational resource, especially since more queries to LLM will be required to customize the ads. We here discuss several plausible cost sharing models that LAS might adopt.
@@
The simplest one will be to charge the advertisers whenever the contents are dynamically modified. For example, the advertiser and LAS commit to a mutual contract on how much the advertiser pay the LAS for each responsive ad, and LAS charges the advertiser whenever it responsively modifies the ad contents. As the advertiser might hope to use responsive ad only if the user context is sufficiently relevant to its ads, the contract might specify when the advertiser’s ads will be responsively changed. Also, LAS may provide several options for advertiser to determine how frequently and largely it should responsively modify the ads.
@@
This model would work well for those charged by cost-per-impression, however, advertisers with cost-per-click or cost-per-conversion may not willingly commit to the contract since they desire to pay the platform only if the actual click or conversion happen, even though the platform indeed responsively modified its ads for every impression. LAS might give several options to the advertisers under which event they will be charged to pay for responsive ad. Ads with cost-per-click model may still want to pay for responsive ads per impression or per click. For the latter, LAS may offer a larger amount of commission per click by predicting the average CTR of the corresponding ads.
@@
LAS must also decide whether or not to account the increased amount of payment for each advertiser as the increment of the actual bid or just regard it as a separate commission. For the former case, the advertisers who are willing to charge more for the responsive ad commission fee will take advantage of getting higher chance of delivering their ads, compared to those who do not use responsive ad option. By doing so, the auction module can select the ad that truly maximize the total revenue expected from delivering a single advertisement, thereby increasing the platform’s revenue.
Finally, we remark that these involved process demand additional computational resources, potentially impacting latency when incorporating dynamic ads. Online advertisement typically requires the overall latency of real time bidding process to be within $100$ms, however, we experience several second of delay to modify the image output presented in Figure 3. Thus, it is technically interesting to balance the tradeoff between the ad’s quality and latency, e.g., naively by using cache, and also it might be possible to pass along part of such cost to the advertisers via monetary charge.

### v3

@@ 4.3 Prediction Module
Prediction module is responsible for computing user’s satisfaction-rate (SR) and click-through-rate (CTR). User’s SR indicates how much the user is satisfied with the given output. This SR affects the final decision making process of LLMA since if there is a chance that the user is very disappointed with the modified output, then it should reconsider displaying that output as it may hurt the user experience a lot. On the other hand, CTR indicates the probability that the user clicks the ad link included within the modified output. This is also critical in determining the auction winner since this directly affects the revenue of the LLMA. More precisely, if an advertiser’s bidding method is cost-per-click (CPC), then CTR proportionally affects the advertisement revenue as the expected revenue of transferring its ad can be computed as CPC times the CTR. Overall, both of SR and CTR can be defined as a function of original output $X$, modified output $X^{\prime}$, query $q$, and context $c$, which returns a real value in $[0,1]$. Detailed methodologies for estimating/learning these functions will be discussed in Section 5.
Comparison to SA
Different from the traditional SA whose output is static (as ad image/hyperlinks), LLMA constructs the output in a highly dynamic manner. This makes the prediction module more difficult to learn the CTR. Moreover, in LLMA, the prediction of user satisfaction is much more directly affected by the incorporation of the ads in the output. This is in stark contrast with SA, whose user experience is usually affected only by the number of ad slots in SERP.
4.4 Auction module
@@
Computed all the required parameters above, we finally run the auction module to determine who wins the auction and how much the advertiser will be charged. The input to the auction module is the tuples of $(\text{bid}_{i},\text{sr}_{i},\text{ctr}_{i})_{i\in[n]}$ which denotes bid amount, satisfaction rate, and click-through-rate, respectively. Output of the auction module is an allocation $a\in\{0,1\}^{n}$ and a payment $p\in\mathbb{R}_{\geq 0}$. Precisely, auction module determines the auction format which consists of the allocation function and the payment function, where the allocation function decides which advertisement to deliver and the payment function computes how much the advertisers pay LLMA. [^11] In its current form, this does not allow the possibility of adjusting the final output in order to balance between multiple advertiser’s preferred output. Our framework, however, does not preclude possibility for such extension, which we will elaborate more in Appendix B.
@@
Importantly, the main goal of LLMA is to maximize its revenue in a long term manner. To this end, this should balance a trade-off between short-term revenue and the user retention. Therefore, the objective can be modeled as a function from bid amount, CTR, and SR to a nonnegative real number that indicates a score of a modified output, i.e., selecting $i^{*}=\operatorname*{argmax}_{i\in[n]}\text{Obj}(\text{sr}_{i},\text{ctr}_{i},\text{bid}_{i})$. We do not discuss how to choose the objective function here, given the vast literature on the sponsored search auction. After designing the score function, one should determine an auction format. Also, given the long line of literature on auction design, one can pick the most preferred mechanism, e.g., first-price auction (FPA) or second-price auction (SPA).
Comparison to SA
The main difference to the SA is, similar to what is discussed in the previous subsection, the user satisfaction is much more important measure to account for. For example, the allocation function in SA represented by (1.1) is social welfare which only accounts for platform and advertisers’ utility. In LLMA, however, one might also need to consider user’s utility as a function of predicted SR, which would change the allocation function of the mechanism and the payment correspondingly.
@@ 5 Challenges
Recall that our overall framework consists of four modules: modification, bidding, prediction, and auction module. For each module, we discuss several characteristics, technical challenges, and corresponding research questions in implementing the proposed architectures in practice, evaluating their overall practicality with respect to the criteria defined in Section 3. We provide further discussions in Appendix B.
@@ 5.1 Modification module
Challenges: Advertiser modification model that guarantees no privacy-issues, reliability, user satisfaction, and low-latency. In the advertiser modification model, LLMA should somehow deliver $q,X,C$ to each $\text{adv}_{i}$. This may leak some private information of the user to the bidders, thereby suffering privacy issues, which is critical for user experience. It should be addressed how LLMA can partially or indirectly disclose the information to prevent user privacy while guaranteeing high quality output from the advertiser modification module.
In addition, the advertiser modification model has reliability issue since there is a possibility that the advertiser-modified output includes some illegal or spamming contents. The LLMA may devise an additional module to maintain its robustness against such kind of adversarial behavior, however, this is still a cost of the LLMA to bring them in the system. This further possibly degrades the user satisfaction since the advertiser-generated output may not be satisfactory for the user compared to what the LLM generates. Furthermore, this requires additional communication between the LLMA and the advertisers, thereby increasing the overall latency. Advertiser modification model faces many concerns as discussed here, thus it might not be practical without an innovative methodology that resolves these issues.
Challenges: Effectively reflecting advertiser preference on the output for LLMA modification model. On the other hand, LLMA modification model does not suffer the same amount of privacy, reliability and latency issues mentioned above. Furthermore, this model would be more in favor of the user since LLMA would aim to improve the user experience. However, one disadvantage is that this may not fully reflect the advertiser’s preference on the output modification process, thereby partly decreasing the advertiser’s satisfaction. Thus, it should be explored how one can effectively reflect advertisers’ preference on the modified output to better capture the output that advertisers would have preferred.
Prospect: Balancing the trade-off between LLMA/advertiser modification models. In order to improve the advertiser’s satisfaction for the LLMA modification model, one approach is to allow the advertiser to submit indirect indicators that represent its preference on what it expects the modified output to be. More in a detail, once the user asks a query $q$, LLMA delivers $q$, the original output $X$, and possibly the context $c$ to the advertiser, and the advertiser returns a string $Y$ that indicates its preference in words. After then, LLMA may apply $Y$ as a prompt in the LLM to generate the modified output. This approach allows more flexibility for the advertiser to express its preference on the output. Note that, however, this still requires additional communication cost compared to the LLMA modification, which possibly increases latency. Another way to better capture advertisers’ preference on valuations without further communication is to ask advertisers to write a list of documents to give it as samples to the LLM as a prompt engineering.
@@ 5.2 Bidding module
Challenges: Implementing dynamic bidding model without privacy and latency issue. The biggest advantage of the dynamic bidding model is that, it would induce higher advertiser’s satisfaction rate, which may be attractive for some advertisers who have technical capability to dynamically send bid with respect to the proposed output. This is because the advertisers have full flexibility in deciding their bids after observing the modified output. Similar to the advertiser modification model, dynamic bidding model may have privacy issue if LLMA discloses some private information to the advertiser. Moreover, dynamic bidding model may suffer additional latency issue since the entire set of modified output should be delivered to the advertisers. Note that there is no reliability concern as only the bid amount is communicated.121212 Note that static bidding model does not suffer any of the privacy, latency, reliability issues. However, as the advertisers cannot fully customize their bid per each query based on the modified output, this may degrade the advertiser satisfaction. Overall, the research in this direction should carefully address how practical the dynamic bidding model is, and how it can be implemented without violating privacy and latency issue. [^13] Further, one may also investigate how the ad market would comprise of, when there is a possibility that the advertisers hire a proxy agent to write bid on behalf of them, and how the proxy agent (or advertisers themselves) can optimize bids in such scenario.
Prospect: Balancing the trade-off between static/dynamic bidding models. In the static bidding model, a natural approach to improve the advertiser’s satisfaction is to enable a more generic static function as the contract between the advertiser and LLM provider. Specifically, LLMA can propose the bidding contract as a function that maps an indirect measure of the modified output to a nonnegative real number representing the bid, i.e., the advertiser’s valuation. For instance, LLMA and an advertiser agrees on a inverse proportional contract based on the similarity distance $d$ between the original output $X$ and the modified output $X_{i}$. For example, if $X_{i}$ is examined to be very different form $X$, then $d(X,X_{i})$ will be large, and then the advertiser’s valuation, i.e., bid, on $X_{i}$ is small since it is believed that the incorporation of ad too much changed the content of the output, thereby hurting user experience and capturing less user attention. Note that in this case, the advertiser may need to partly understand how LLMA estimates the distance measure and how it’s defined. More generally, LLMA might construct a more refined way to measure the user’s interest, attention, relevance for the modified output from the perspective of $\text{adv}_{i}$, and let the advertisers commit to either of the contract based on its taste. These are crucial aspects to explore to run LLMA in practice.
@@ 5.3 Prediction module
Challenge: Efficient and precise implementation of prediction module. The process of estimating CTR will be analogous to that in the modern online advertisement system. We can train a prediction system to predict $\text{ctr}_{i}\in[0,1]$ given the input $X_{i},q,c$ for each $i\in[n]$ based on the feedback data from the user. [^14] For instance, as is typical in the modern online advertisement system, one may use factorization machine based online algorithm after extracting some features from user, context, and the queries. More detailed discussion can be found in McMahan et al. (2013). As discussed in Section 4.3, since $X_{i}$ will be very sparsely appear if the entire document is considered as an embedding, we may extract some useful features from each $q,X_{i}$ and $c$, and learn the mapping from features to the true value. Whether the user regenerates the response, clicks the ad, or exits the LLM, can be used as feedback data to improve the prediction module.
@@
Challenge: Relevance/similarity distance measure to estimate user satisfaction. To estimate SR, one may first consider an explicit approach given an assumption that the original output is optimal. If we assume that the original output is optimal, then essentially a distance between the original output $X$ and the modified output $X_{i}$ will be a good measure to predict the SR since if $X_{i}=X$, then the user will be the most satisfied. In this context, given the original $X$, one may use the marginal probability ${\mathrm{Pr}\left[\,X|q\,\right]}$ as an indicator for SR. For example, one can define output distance by $d(X,X_{i}):=||{\mathrm{Pr}\left[\,X|q\,\right]},{\mathrm{Pr}\left[\,X_{i}|q\,\right]}||$ for some proper norm. Note that the probability for an output to be obtained can be computed via standard method in the literature (Vaswani et al., 2017), by computing the probability distribution in a token-by-token manner. Further, one can consider more general distance function $d$, e.g., the semantic similarity of two documents/phrases (Mikolov et al., 2013; Cer et al., 2018; Conneau et al., 2019). Essentially, a concrete research question in this context is to identify a good measure to capture the user’s satisfaction in advance when ads are incorporated in the output.
@@
Prospect: Incorporating distance measure and online learning. As discussed above, one may consider combining the approaches of distance measure and online learning from user feedback for SR prediction. To online learn SR estimates from user feedback, a useful indicator of whether the user is satisfied with the output is whether the user re-generates the output given the same (or similar) query. Hence, one may aim to learn a function which outputs the sr, given the query $q$, modified output $X^{\prime}$, and context $c$. This approach does not assume that the original output $X$ is indeed optimal, thereby allowing the possibility that the user may be satisfied with a modified output $X_{i}$ even though its distance from $X$ is measured to be far. Note that one should implement an independent training system to operate this learning process, leading to additional burden for LLMA. To incorporate advantages of these two approaches, one can consider an online learning model where the distance is also integrated as one of the features to predict it, i.e., $d(X,X_{i})$ is used as a feature of the SR prediction system. In case $d(X,X_{i})$ indeed has some positive correlations with the true user satisfaction rate, this would increase the accuracy of the prediction. Essentially, an effective way to capture the both advantages of online learning and distance measure should be studied thoroughly.
@@ 5.4 Auction module
Challenge: Incorporating multiple ads in a single output. Recall that in our auction module, the modified output is fully determined based on the allocation function so that if it is allocated to $\text{adv}_{i}$, then the modified output $X_{i}$ will be displayed to the user. One approach to generalize this framework is, to let the final output to be displayed the one that does not necessarily belong to $\{X_{i}\}_{i\in[n]}$, but rather a new output $X^{\prime}$ that somehow interpolates $\{X_{i}\}_{i\in[n]}$. [^15] We discuss further comparison to the token auction model by Duetting et al. (2023) in Appendix B.1. Let us denote the new output $X^{\prime}$ by balanced output, e.g., see Table 2 in Duetting et al. (2023). By doing so, it might be possible to deliver multiple advertisements in a fair manner, [^16] Another way to display multiple ads at once will is discussed in Appendix B.2 under our original framework. thereby allowing LLMA to bring more revenue by charging multiple advertisers at once. One subtle issue is that, since each advertiser bids $b_{i}$ on delivering $X_{i}$, they may not want to write the same bid for the balanced output $X^{\prime}$, thus it may degrade the advertiser’s experience on participating in the LLMA. [^17] In the static bidding model, as discussed in Section 5.2, this might be handled by committing to a contract based on measures that represents the advertiser’s preferences more in a refined manner. In the dynamic bidding model, one approach would be to append an additional step of asking for bids for the final output again to the advertisers. In summary, incorporating multiple ads in a single response is technically challenging, which would require both theoretical and practical studies to validate reasonable approaches.
@@ 6 Dynamic Creative Optimization via LLM
Figure 4: Integrating responsive advertisement in the images generated by ChatGPT 4. In the left panel, a user requests car recommendations within the price range of \$40,000 to \$60,000, and the chat bot, when capable of advertising, generates a “BMW 3 Series” image (left). Personalized ads modify this image for a 60-year-old woman (middle) or man (right) based on user context like age and gender. In the right panel, a user seeks machine learning conferences in Florida. The chat bot advertises “SunShine Travel” with its logo on a generic Florida image (left). With personalized ads, the model tailors promotions, offering flight tickets from the user’s location and showing more relevant Florida images. For example, for a young unmarried person from DC, the output advertises DC to Florida flights (middle) and features an image of beach there, while for a married person in California, it promotes California to Florida flights with images of amusement parks. This enhances user experience with targeted and appealing content.
@@
Now, we discuss about the possibility of advanced dynamic creative optimization by using LLM. Dynamic creative optimization (DCO)  [^18] This is often called as dynamic product ads (DPA), but DPA typically refers to the process of selecting which item to advertise from a given catalog of an advertiser, whereas DCO denotes the process of modifying the ad itself., in the traditional search and display ad markets, refers to the technique of dynamically adjusting the combination of ad assets and constructing an advertisement that best fits the customer’s preference. It can markedly enhance the quality and relevance of the ad to the user by tailoring the contents of ad in a personalized manner. Many online advertisement platforms adopt DCO-based ads, which is often called as responsive ad (GoogleDisplayAd, ; GoogleSearchAd, ).  [^19] As traditional DCO is restricted to dynamically construct the complete image file of ad, we here restrict our attention to the image-based ads as well. However, it is apparent that LLM based DCO would work beyond image-based ads.
@@
In the canonical DCO framework, a static ad consists of a single image file that can only be displayed in a specific set of ad channels, whereas dynamically optimized ad consists of a single ad template and multiple options for each asset that fills in the ad template. This is typically done by the advertiser registering multiple options for each asset given an ad template, and then the system picks the optimal combination. This procedure aims at increasing the CTR and SR, which would eventually improve the welfare efficiency of the online ad ecosystem (McMahan et al., 2013) as well as the revenue of the ad platform.
@@
Unlike from the traditional DCO techniques, the arise of LLM, especially those support the output with image such as ChatGPT 4, enables more involved process of dynamic creation of ads, as well as replaces the role of traditional DCO in online ad platforms. For example, given a user query and context, the platform may send a query to LLM to customize the image of ad to attract the user by capturing user preference. We provide representative examples in Figure 4 where two scenarios of personalizing ads is depicted. Note that the use of LLM-based DCO is not restricted to LLMA, but can be broadly applied to SA and display ad.
@@
Challenge: Refining each module. To integrate responsive ads, a more advanced modification module is imperative. This module should not only enable the alteration of the original response to include ads but also incorporate them in a manner that takes into account user preferences. This entails utilizing user contexts, e.g., gender/location/the current device used for the query, in the ad generation process. To be more specific, the modification module should incorporate user context in the process of modifying the original response. This makes ads more appealing to the user by considering their preferences. Leveraging language models, we can prompt them to factor in various elements related to user context during the generation of modified outputs. As illustrated in Figure 4, information about users’ context can be utilized to create more captivating advertisements for them.
Further, the prediction module may interact with the modification module to commit to the modified output that mostly increases user’s experience, i.e., CTR and SR. Even for SA, the prediction module may need to be made more sophisticated to predict the CTR and SR of the responsive ads. This is in stark contrast to the standard DCO since it may not significantly change the ad image overall, but only slightly changes the part of it by replacing some ad assets. Correspondingly, the advertiser may value the generated ad differently, i.e., write different bid, with respect to the quality of the ad generated by LLM-based DCO. Overall, a vast amount of studies should be done to clearly investigate a plausible framework for LLMA, even for SA, to adapt the LLM-based DCO.
Challenge: Meeting system requirements. Obviously, running LLM-based DCO should not entail the failure of the system requirements discussed in Section 3. For instance, in terms of latency, we observed that generating an image with the GPT4 takes almost about $10$s for Figure 4. Given that the online advertising typically requires the ad latency up to $100$ms, the required latency is almost $100$ times the conventional threshold. Thus, the fundamental research challenge is to enable a fast production of an ad image to meet the conventional latency requirement in the online advertising system. Further, it will be technically interesting to balance the tradeoff between the ad’s quality and latency which governs the advertiser and user utility, respectively.
Interestingly, as acute readers may have noticed, Figure 4 includes some technical issues in the images themselves. For example, the first image in the car ads generate a word that is not easily understandable. Indeed, LLM-based DCO may induce the reliability issue due to the hallucinating nature of LLM (Ji et al., 2023; Guerreiro et al., 2023; Bubeck et al., 2023). Thus, LLM-based DCO should ensure that it does not generate a hallucinated ad that may hurt the advertiser satisfaction. A simple approach might be to require LLM to minimally revise the reference images, requesting the advertisers to submit several images for reference.
@@
Challenge: Cost sharing model. Unlike from the traditional DCO technique that simply determines the efficient combination of ad assets, LLM-based responsive ad can do much beyond by creating truly new content. On the other hand, such process necessarily entails increased use of computational resource, especially since more queries to LLM will be required to customize the ads. We here discuss several plausible cost sharing models that LLMA might adopt.
@@
The simplest one will be to charge the advertisers whenever the contents are dynamically modified. For example, the advertiser and LLMA commit to a mutual contract on how much the advertiser pay the LLMA for each responsive ad, and LLMA charges the advertiser whenever it responsively modifies the ad contents. As the advertiser might hope to use responsive ad only if the user context is sufficiently relevant to its ads, the contract might specify when the advertiser’s ads will be responsively changed. Also, LLMA may provide several options for advertiser to determine how frequently and largely it should responsively modify the ads.
@@
This model would work well for those charged by cost-per-impression, however, advertisers with cost-per-click or cost-per-conversion may not willingly commit to the contract since they desire to pay the platform only if the actual click or conversion happen, even though the platform indeed responsively modified its ads for every impression. LLMA might give several options to the advertisers under which event they will be charged to pay for responsive ad. Ads with cost-per-click model may still want to pay for responsive ads per impression or per click. For the latter, LLMA may offer a larger amount of commission per click by predicting the average CTR of the ads.
@@
LLMA must also decide whether or not to account the increased amount of payment for each advertiser as the increment of the actual bid or just regard it as a separate commission. For the former case, the advertisers who are willing to charge more for the responsive ad commission fee will take advantage of getting higher chance of delivering their ads, compared to those who do not use responsive ad option. By doing so, the auction module can select the ad that truly maximize the total revenue expected from delivering a single advertisement, thereby increasing the platform’s revenue.